OpenStudy (anonymous):

How many solutions does the equation 3x − 7 = 4 + 6 + 4x have? Two None Infinitely many One

Elsa213 (elsa213):

the first equation can be re-arranged 3x – 4y = 12 3x = 12 + 4y 4y = 3x - 12 y = 3/4 -12

Elsa213 (elsa213):

so we have these two equations: y = 3/4x -12 y = 4x – 3

Elsa213 (elsa213):

so I can just solve the system instead :) 3x – 4y = 12 ,plug y=4x-3 3x - 4(4x-3) = 12 3x -16x +12 = 12 3x-16x = 0 -13 x = 0 x = 0.....

Elsa213 (elsa213):

y = 4x – 3 y = -3 3x – 4y = 12 -4y = 12 y = -3 there seems to be a solution

OpenStudy (anonymous):

so theres two solutions? @Elsa213

Elsa213 (elsa213):

yes c: x=0 and y=-3